OpenStudy (anonymous):

Solve the following system of equations. 3x - 2y = 6 6x - 4y = 14 (0, 0) (6, 14) Infinitely Many Solutions No Solutions

hero (hero):

3x - 2y = 6 6x - 4y = 14 y = (6 - 3x)/-2 y = (14 - 6x)/-4 y = y -4(6 - 3x) = -2(14 - 6x) -24 +12x = -28 +12 x -24 = -28 False

hero (hero):

No solution